In 1992, the FAA conducted 86,991 pre-employment drug tests on job applicants who were to be engaged in safety and security-related jobs, and found that 1,143 were positive. (a) Construct a 95 percent confidence interval for the population proportion of positive drug tests. (b) Why is the normality assumption not a problem, despite the very small value of p?
Answer by stanbon(75887)

In 1992, the FAA conducted 86,991 pre-employment drug tests on job applicants who were to be engaged in safety and security-related jobs, and found that 1,143 were positive. (a) Construct a 95 percent confidence interval for the population proportion of positive drug tests. (b) Why is the normality assumption not a problem, despite the very small value of p?
------------------------------
p-hat = 1143/86991 = 0.01316...
E = 1.96*sqrt[0.01316*0.98684../86991] = 0.0003863
--------------
95% CI: 0.01316-0.0003863 < p < 0.01316+0.0003863
====================================================
n is so large np > 5 and nq > 5.
Cheers,
Stan H.